OpenStudy (anonymous):

What is the solution set of (x - 2)(x - 3) = 2

OpenStudy (anonymous):

expand first \[x^2-3x-2x+6=2\]\[x^2-5x+6=2\]

OpenStudy (anonymous):

\[x^2-5x+6-2=0\]\[x^2-5x+4=0\]solve using the quadratic formula\[-b \pm \sqrt{b^2-4*a*c}\]--------------------\[ 2a\]

myininaya (myininaya):

(x-4)(x-1)=0 x=4 x=1